# The Physics and Metaphysics of Social Powers: Bridging Cognitive Processing and Social Dynamics, a New Perspective on Power through Active Inference

### v1

@@ 2.2 Narratives as Scripts
The role of narratives in power dynamics, as scripts which normalize behavioral patterns by providing templates for “appropriate” action, may make “deviant” behavior appear explorative (others may learn, inspired by their curiosity) or inappropriate (Constant et al., 2018): leading to social effects such as castigation or neglect. This leads to the creation of shared meaning as well as to the justification of existing power structures. Because of the challenges inherent to the coordination of social reality—as mentioned: whether agents following a script are (critically) aware of the implied norm, or simply perform it, being the main question—we can also think about scripting narratives in terms of self-fulfilling prophecies, hype, fictions that make themselves real, etc.: none of the possibilistic options on social offer need be realistic, so long as they are strong enough to attract attention (driven by the desire to reduce uncertainty about future states).
@@
The effects of variational free energy-minimization are observed at the level of the basic structures of language: (semantic) attractors such as redundancy, rhyme, alliteration, etc., can be understood as patterns which aid memory, communication and information retrieval policies, because they permanently reappear, pooling attention to the “same” phenomena, thus greatly reducing the need to interpret and organize linguistic interaction and cognition anew. We can understand, for example, the possible emergence of rhythmic patterning in communication as enabling the storage and access of information collectively: by connecting a sound-pattern to a phenomenon, we reduce uncertainty about reality once that sound-pattern is witnessed, in true Pavlovian fashion. By way of the same uncertainty-minimizing mechanisms, (simple or complex) narrative constructs create cognitive situations where agents reduce complex aspects of the world down to preferred scenarios: it is easier to imagine the end of the world than the end of capitalism (Jameson/Žižek in Fisher, 2009). In light of our arguments: this might be, perhaps, because the end of the world is relatively easy to imagine: asteroid collision, megavolcano, etc., while the systemic nature of capitalism is rather difficult to pull apart in detail. Or: it is easier to reduce uncertainty about my current stress-predicaments by blaming my neighbor’s lifestyle, than by reassessing my entire lifestyle (generative model) and change it on the basis of a myriad other variables. Cognition, often led by a desire for and mechanics of the shortest path, tends towards predictable, digestible narratives more often than not.
@@
Unlike short-range projective experiences such as the ones detailed in Maisto et al., 2024, where prediction errors present immediate challenges, narratives and scripts provide vastly speculative semantic spaces for exploring models against hypothetical outcomes. As we have been highlighting, with regard to power, the problems emerge in situations where agents controlling dominant narratives are guided by nothing other than power-acquisition or maintenance, leading not only to, e.g., misinformation, echo chambers, conspiracy theories, etc. (Albarracin et al., 2022), but, in political communication and organization, to the justification of future policies based on (supposedly) given conditions (deriving an ought from an is, essentially). Abstract semantic constructs (fictions, narratives, scripts), which can be understood as constraint-regimes coming to operate as distributed generative models, create cognitive situations where agents can learn by “internally” simulating responses to high-stakes scenarios, while maintaining corporeal integrity and safety, all the while plastically exploring their epistemic condition. An environment pooling attention to strong-enough narratives will elicit the scripting of policies in line with those narratives (as mentioned, regardless of their ‘veridical’ content). For agents under thermodynamic duress (socioeconomic, corporeal, etc.): the simpler the narrative, often, the better. This being for the banal reason that they are under duress, and therefore incapable of making enough spacetime available for deep, projective computations. Narrative structures tempered by the pooling of attention act as semantic attractors which organize possible world-states into explorable patterns. These patterns function as probability-possibility landscapes for belief-updating, allowing agents to or restricting them from fine-tuning their predictive models through engagement. Considering homo sapiens’ universal tendency towards narrative-sharing, this process can be understood as a corollary of actively-inferential social agents, particularly if we understand ‘control’ as the phenomenon of conscious attention to how the modulation of variables effectuate the production of the future151515 In the context of “fictional inference” (de Jager 2023), which is much of what humans do, this is the exploration of domains where direct experience would be prohibitively risky or even impossible. Fiction provides “all of the learning, at relatively low risk” (ibid.).. The theoretical framework developed here has significant implications for understanding both the cognitive foundations of narrative engagement and how the power of attention to these results in the scripting of communal behavior. We now move to an analysis of power through the lens of AIF.
@@ 3.1 Individual agency and empowerment
Autonomy refers to a system’s capacity to govern itself based on internally defined norms (Watson and Scheidt, 2005). Autonomous behavior emerges through entropy maximization under constraints. Agents do not simply maximize randomness (though in some instances they do: explorative behavior is subject to high noise and benefits from high noise, again, note on average long-term VFEM in social agents). Agents optimize their behavior by maintaining a balance between maximizing entropy and respecting system constraints (Jaynes, 1957). This is consistent with Jaynes’ principle of maximum entropy, which states that the least biased estimate of a probability distribution is the one that maximizes entropy subject to known constraints. In autonomous agents, these constraints include both the agent’s internal model and environmental dynamics (Golan and Foley, 2022).
We can think of empowerment as a formal measure of an agent’s control over its future states, defined as the maximum mutual information between an agent’s actions and the resulting states (Klyubin et al., 2005):
@@
where $I(A_{t,n};s_{t+n})$ denotes the mutual information between the next $n$ actions $A_{t,n}$ from time $t$ and the state $s_{t+n}$ at time $t+n$, and $p(a_{t,n})$ represents the probability distribution over the next $n$ actions from time $t$ (Hynes, 2024). In this context, agents select actions that maximize their control over future environmental states, optimizing their ability to influence outcomes. Higher empowerment indicates greater potential for an agent to achieve desired future states. Importantly, empowerment can be decomposed into entropy terms as:
@@
where $H[X]$ represents the Shannon entropy of the random variable $X$ and $H[X,Y]$ denotes the joint entropy of $X$ and $Y$. This decomposition reveals that empowerment involves maximizing the marginal entropies of state and action distributions while maintaining strong correlations between actions and their consequences (Hynes, et al., 2024). We can also link these ideas to recent developments in understanding intrinsic motivation in dynamical control systems (Tomkin et al., 2024) and complex behavior emerging from the drive to occupy future action-state path spaces (Ramírez-Ruiz et al., 2024).
Empowerment as defined above is an intrinsic motivation that does not specify how an agent should take actions to achieve specific goals. Extensions have been introduced to incorporate goal-directedness by introducing a reward constraint (Volpi and Polani, 2020) or tradeoff (Leibfried et al., 2019), which highlight the tensions that may arise between intrinsic and extrinsic motivators.
Under the Active Inference Framework (AIF), autonomous agents select policies (sequences of actions) by minimizing expected free energy, which includes both epistemic (information-seeking) and pragmatic (goal-directed) terms (Friston et al., 2010; Parr et al., 2022). AIF differs from empowerment in several respects, but one crucial point of difference is that it explicitly models the partial observability inherent in many decision-making environments, motivating agents to resolve uncertainty through curiosity-driven behavior. This balance between exploration and exploitation allows agents to both gather information about their environment and pursue specific goals. Policy selection in autonomous agents involves maintaining a generative model of environmental dynamics. To do so, agents must update their beliefs about current states through perceptual inference. Then, they can select actions that minimize expected free energy over future states and ultimately learn from action outcomes to improve the generative model.
@@ 3.2 Social power dynamics
Powerful entities offer perceived stability and predictability in an uncertain environment. In AIF terms, they help other agents minimize their expected free energy by providing reliable frameworks for prediction, essentially by dictating the information geometry landscape (Constant et al., 2019). Power confers significant advantages in information processing and management. Powerful entities occupy positions that grant them privileged access to information streams, enabling more accurate modeling of their environment (Friston et al., 2020). Ultimately, they also control what can and cannot be modeled, and in what way, through their position and resources, powerful agents can offload computational demands onto others or their environment, allowing for more efficient processing of complex information. In the complexifications of the “free market” condition unfolding over the 20th century, we can understand these dynamics as the emergence of a novelty-generating network of agents which has proven to provide benefits to those with leverage over its structures: agents already well-positioned within the informatic domain. In sociological lingo: capitalist capture, or the cancellation of the future (Fisher, 2009). Power influences how errors and uncertainty are processed within social systems. Powerful entities can better absorb or reinterpret errors due to their broader resource base and influence over narrative construction. Through their ability to shape both physical and informational environments, powerful agents can actively reduce uncertainty for themselves while potentially increasing it for others.
@@ 3.3 Formal Framework Integration
Power dynamics within and between agents can be formalized by addressing various dimensions of control, influence, and synchronization. At the individual level, recall that empowerment quantifies an agent’s potential influence over its environment by assessing the mutual information between the agent’s actions and the resulting states (Klyubin et al., 2005). Here, we discuss some implications of this notion in the context of active inference.
@@
where $G(\pi)$ represents expected free energy for a policy $\pi$, and $D(\pi,o)$ is the deontic value associated with aligning actions with social norms and expectations (Constant et al., 2019). Social norms modulate policy selection, encouraging behavior that is goal-oriented and aligned with collective expectations. This deontic value may also be linked to a particular kind of empowerment of other agents. Consider that to another agent, one’s own actions may be viewed as part of the external state that is being predicted and controlled. Therefore, the empowerment of other agents includes their influence on one’s own policy, which may be reflected by the deontic value term above. A deeper investigation into how the two are precisely related is beyond the scope of this paper, but would shed light on the relationship between social norms and individual/collective empowerment.
@@
where $Q_{i}(s)$ and $Q_{j}(s)$ denote the belief distributions of agents $i$ and $j$ over the hidden state, respectively (Friston et al., 2020). This quantity reflects the degree of belief alignment, indicating the extent to which a shared understanding emerges from social interaction. The ways in which this is unconscious, tacit, or otherwise unexplored are the elements which lead to many of the intractable aspects of formalizing these dynamics. This is a consideration which ought to be explicitly stated in any modeling attempt.
@@
Linking this idea back to empowerment, notice that defining “external states” is more complicated in social contexts, because one agent’s internal state is another agent’s external state, and vice versa. This means that different agents’ internal representations of the world must be similar in some parts (those that are external to all of them) and different in other parts. Therefore, part of what it means for an agent to maximize their empowerment in a social context is to actively make parts of other agents more predictable and useful to themselves, which is often more easily achieved by bringing other agents’ generative models in line with one’s own. In general, belief alignment can emerge as a natural consequence of each agent’s drive to minimize free energy within the right social context. This fosters synchronized behavior, shared narratives, and collective understanding, unifying individual agency, social influence, and collective behavior under the cohesive formalism of AIF (Friston et al., 2020). However, as the brief analysis above reveals, such synchronization through belief alignment may come about as a result of the exertion of social power, highlighting the importance of understanding not just when this synchronization occurs, but also how it does so. As we conceptualized it, power is an agent’s capacity to shape and control the (semantic) “attractors” within a social or informational landscape. These attractors represent stable states or patterns that agents are drawn towards, which reduce uncertainty and create predictable social structures. Power enables agents to influence which states become salient or preferred for others, thereby controlling the flow of attention, shaping beliefs, and modulating the precision of predictions. The 20th and 21st century have shown, time and time again—especially so in the last 20 years—how mediatic domination (narrative control) results in the shaping of attractors which organize life.
@@
In leader-follower dynamics, a leader minimizes this divergence by exerting control over precision, amplifying others’ confidence in specific interpretations of their sensory data. For example, if the leader’s goal is clearly defined and they select social epistemic actions (more informative, albeit less efficient actions), they convey their intentions early and reduce uncertainty for followers (Maisto, Donnarumma, & Pezzulo, 2024). This approach encourages other agents to update their beliefs to align with the leader’s, effectively lowering KL divergence across the group. Even in cases where leadership is not predefined, an agent can assume a leadership role by adjusting precision to show high confidence in their beliefs or goals. This increase in precision draws other agents toward the confident agent’s stance, resulting in a convergence on shared beliefs. Precision-driven leadership can thus promote belief alignment without coercion, as agents gradually follow those who express strong, reliable predictions. In this way, the power to direct attention, structure narratives, and therefore modulate precision enables agents to actively shape collective beliefs, creating cohesion within groups and guiding group behavior through the minimization of belief divergence.
@@
The applications of this inevitably begin at reductive formalizations (and the math is never the territory (Andrews, 2021)), however, the implications of these arguments for complex social dynamics in the realms of, e.g., policy and governance are vast. Coordination dynamics at the level of simple perception-action cycles such as the ones analyzed in Maisto et al. 2024, need to be framed in terms of their possibilities for novel understandings of social change processes: how should humanity—as one, as many—develop strategies with enough coherence and plasticity so as to resist the oppressive aspects of dominant narratives which do not benefit the coordination of the whole? What kinds of timeframes are we thinking about (in terms of generations, or the reorganization of our understanding of future time altogether)? Leader-based top-down control ‘works’ when established cooperation dynamics are already in place, and the same can be said about distributed control (either as presented by a socially-positive cybernetics (e.g., Beer) or “free market”, invisible hand narratives).
@@
A starting point for the analysis of such multi-agent interactions may be from the point of view of game theory. In particular, the concept of Free-Energy Equilibria (FEE) captures the notion of a joint policy in which all agents in a given interaction are minimizing their (Expected) Free Energy (Hyland et al., 2024). Under certain conditions, it may be expected that interactions between agents tend to drive the system toward FEE. However, as is well understood in game theory, there are typically many equilibria of a game and depending on the starting point and interaction/learning dynamics, the system may converge to undesirable equilibria where a handful of agents exert significant control over others to their detriment. An important area of future study is thus to first establish a better understanding of what makes different equilibria more or less socially desirable, and then to understand how incentive landscapes may be shaped to steer societal outcomes toward better equilibria. Shared narratives, where agents are in the know about what is shared, are fundamental here, and this begins by paying attention to the (possible) rules in place, hence the need for formal methods. This is part of what critical metacognition entails: knowing how we are collectively navigating uncertainty, and taking account of already-given asymmetries. In the conclusion below, we explore a few salient implications of our arguments.

### v2

@@ 2.2 Narratives as Scripts
The role of narratives in power dynamics, as scripts which normalize behavioral patterns by providing templates for “appropriate” action, may make “deviant” behavior appear explorative (others may learn, inspired by their curiosity) or inappropriate [Constant et al., 2019]: leading to social effects such as castigation or neglect. This leads to the creation of shared meaning as well as to the justification of existing power structures. Because of the challenges inherent to the coordination of social reality—as mentioned: whether agents following a script are (critically) aware of the implied norm, or simply perform it, being the main question—we can also think about scripting narratives in terms of self-fulfilling prophecies, hype, fictions that make themselves real, etc.: none of the possibilistic options on social offer need be realistic, so long as they are strong enough to attract attention (driven by the desire to reduce uncertainty about future states).
@@
The effects of variational free energy-minimization are observed at the level of the basic structures of language: (semantic) attractors such as redundancy, rhyme, alliteration, etc., can be understood as patterns which aid memory, communication and information retrieval policies, because they permanently reappear, pooling attention to the “same” phenomena, thus greatly reducing the need to interpret and organize linguistic interaction and cognition anew. We can understand, for example, the possible emergence of rhythmic patterning in communication as enabling the storage and access of information collectively: by connecting a sound-pattern to a phenomenon, we reduce uncertainty about reality once that sound-pattern is witnessed, in true Pavlovian fashion. By way of the same uncertainty-minimizing mechanisms, (simple or complex) narrative constructs create cognitive situations where agents reduce complex aspects of the world down to preferred scenarios: it is easier to imagine the end of the world than the end of capitalism [Fisher, 2009]. In light of our arguments: this might be, perhaps, because the end of the world is relatively easy to imagine: asteroid collision, megavolcano, etc., while the systemic nature of capitalism is rather difficult to pull apart in detail. Or: it is easier to reduce uncertainty about my current stress-predicaments by blaming my neighbor’s lifestyle, than by reassessing my entire lifestyle (generative model) and change it on the basis of a myriad other variables. Cognition, often led by a desire for and mechanics of the shortest path, tends towards predictable, digestible narratives more often than not.
@@
Unlike short-range projective experiences such as the ones detailed in [Maisto et al., 2024], where prediction errors present immediate challenges, narratives and scripts provide vastly speculative semantic spaces for exploring models against hypothetical outcomes. As we have been highlighting, with regard to power, the problems emerge in situations where agents controlling dominant narratives are guided by nothing other than power-acquisition or maintenance, leading not only to, e.g., misinformation, echo chambers, conspiracy theories, etc. [Albarracin et al., 2022], but, in political communication and organization, to the justification of future policies based on (supposedly) given conditions (deriving an ought from an is, essentially). Abstract semantic constructs (fictions, narratives, scripts), which can be understood as constraint-regimes coming to operate as distributed generative models, create cognitive situations where agents can learn by “internally” simulating responses to high-stakes scenarios, while maintaining corporeal integrity and safety, all the while plastically exploring their epistemic condition. An environment pooling attention to strong-enough narratives will elicit the scripting of policies in line with those narratives (as mentioned, regardless of their ‘veridical’ content). For agents under thermodynamic duress (socioeconomic, corporeal, etc.): the simpler the narrative, often, the better. This being for the banal reason that they are under duress, and therefore incapable of making enough spacetime available for deep, projective computations. Narrative structures tempered by the pooling of attention act as semantic attractors which organize possible world-states into explorable patterns. These patterns function as probability-possibility landscapes for belief-updating, allowing agents to or restricting them from fine-tuning their predictive models through engagement. Considering homo sapiens’ universal tendency towards narrative-sharing, this process can be understood as a corollary of actively-inferential social agents, particularly if we understand ‘control’ as the phenomenon of conscious attention to how the modulation of variables effectuate the production of the future151515 In the context of “fictional inference” [de Jager, 2023], which is much of what humans do, this is the exploration of domains where direct experience would be prohibitively risky or even impossible. Fiction provides “all of the learning, at relatively low risk” (ibid.).. The theoretical framework developed here has significant implications for understanding both the cognitive foundations of narrative engagement and how the power of attention to these results in the scripting of communal behavior. We now move to an analysis of power through the lens of AIF.
@@ 3.1 Individual agency and empowerment
Autonomy refers to a system’s capacity to govern itself based on internally defined norms [Watson and Scheidt, 2005]. Autonomous behavior emerges through entropy maximization under constraints. Agents do not simply maximize randomness (though in some instances they do: explorative behavior is subject to high noise and benefits from high noise, again, note on average long-term VFEM in social agents). Agents optimize their behavior by maintaining a balance between maximizing entropy and respecting system constraints [Jaynes, 1957]. This is consistent with Jaynes’ principle of maximum entropy, which states that the least biased estimate of a probability distribution is the one that maximizes entropy subject to known constraints. In autonomous agents, these constraints include both the agent’s internal model and environmental dynamics [Golan and Foley, 2022].
We can think of empowerment as a formal measure of an agent’s control over its future states, defined as the maximum mutual information between an agent’s actions and the resulting states [Klyubin et al., 2005]:
@@
where $I(A_{t,n};s_{t+n})$ denotes the mutual information between the next $n$ actions $A_{t,n}$ from time $t$ and the state $s_{t+n}$ at time $t+n$, and $p(a_{t,n})$ represents the probability distribution over the next $n$ actions from time $t$ [Klyubin et al., 2005]. In this context, agents select actions that maximize their control over future environmental states, optimizing their ability to influence outcomes. Higher empowerment indicates greater potential for an agent to achieve desired future states. Importantly, empowerment can be decomposed into entropy terms as:
@@
where $H[X]$ represents the Shannon entropy of the random variable $X$ and $H[X,Y]$ denotes the joint entropy of $X$ and $Y$. This decomposition reveals that empowerment involves maximizing the marginal entropies of state and action distributions while maintaining strong correlations between actions and their consequences. We can also link these ideas to recent developments in understanding intrinsic motivation in dynamical control systems [Tiomkin et al., 2024] and complex behavior emerging from the drive to occupy future action-state path spaces [Ramírez-Ruiz et al., 2024].
Empowerment as defined above is an intrinsic motivation that does not specify how an agent should take actions to achieve specific goals. Extensions have been introduced to incorporate goal-directedness by introducing a reward constraint [Volpi and Polani, 2020] or tradeoff [Leibfried et al., 2019], which highlight the tensions that may arise between intrinsic and extrinsic motivators.
Under the Active Inference Framework (AIF), autonomous agents select policies (sequences of actions) by minimizing expected free energy, which includes both epistemic (information-seeking) and pragmatic (goal-directed) terms [Friston et al., 2010, Parr et al., 2022]. AIF differs from empowerment in several respects, but one crucial point of difference is that it explicitly models the partial observability inherent in many decision-making environments, motivating agents to resolve uncertainty through curiosity-driven behavior. This balance between exploration and exploitation allows agents to both gather information about their environment and pursue specific goals. Policy selection in autonomous agents involves maintaining a generative model of environmental dynamics. To do so, agents must update their beliefs about current states through perceptual inference. Then, they can select actions that minimize expected free energy over future states and ultimately learn from action outcomes to improve the generative model.
@@ 3.2 Social power dynamics
Powerful entities offer perceived stability and predictability in an uncertain environment. In AIF terms, they help other agents minimize their expected free energy by providing reliable frameworks for prediction, essentially by dictating the information geometry landscape [Constant et al., 2019]. Power confers significant advantages in information processing and management. Powerful entities occupy positions that grant them privileged access to information streams, enabling more accurate modeling of their environment. Ultimately, they also control what can and cannot be modeled, and in what way, through their position and resources, powerful agents can offload computational demands onto others or their environment, allowing for more efficient processing of complex information. In the complexifications of the “free market” condition unfolding over the 20th century, we can understand these dynamics as the emergence of a novelty-generating network of agents which has proven to provide benefits to those with leverage over its structures: agents already well-positioned within the informatic domain. In sociological lingo: capitalist capture, or the cancellation of the future [Fisher, 2009]. Power influences how errors and uncertainty are processed within social systems. Powerful entities can better absorb or reinterpret errors due to their broader resource base and influence over narrative construction. Through their ability to shape both physical and informational environments, powerful agents can actively reduce uncertainty for themselves while potentially increasing it for others.
@@ 3.3 Formal Framework Integration
Power dynamics within and between agents can be formalized by addressing various dimensions of control, influence, and synchronization. At the individual level, recall that empowerment quantifies an agent’s potential influence over its environment by assessing the mutual information between the agent’s actions and the resulting states [Klyubin et al., 2005]. Here, we discuss some implications of this notion in the context of active inference.
@@
where $G(\pi)$ represents expected free energy for a policy $\pi$, and $D(\pi,o)$ is the deontic value associated with aligning actions with social norms and expectations [Constant et al., 2019]. Social norms modulate policy selection, encouraging behavior that is goal-oriented and aligned with collective expectations. This deontic value may also be linked to a particular kind of empowerment of other agents. Consider that to another agent, one’s own actions may be viewed as part of the external state that is being predicted and controlled. Therefore, the empowerment of other agents includes their influence on one’s own policy, which may be reflected by the deontic value term above. A deeper investigation into how the two are precisely related is beyond the scope of this paper, but would shed light on the relationship between social norms and individual/collective empowerment.
@@
where $Q_{i}(s)$ and $Q_{j}(s)$ denote the belief distributions of agents $i$ and $j$ over the hidden state, respectively. This quantity reflects the degree of belief alignment, indicating the extent to which a shared understanding emerges from social interaction. The ways in which this is unconscious, tacit, or otherwise unexplored are the elements which lead to many of the intractable aspects of formalizing these dynamics. This is a consideration which ought to be explicitly stated in any modeling attempt.
@@
Linking this idea back to empowerment, notice that defining “external states” is more complicated in social contexts, because one agent’s internal state is another agent’s external state, and vice versa. This means that different agents’ internal representations of the world must be similar in some parts (those that are external to all of them) and different in other parts. Therefore, part of what it means for an agent to maximize their empowerment in a social context is to actively make parts of other agents more predictable and useful to themselves, which is often more easily achieved by bringing other agents’ generative models in line with one’s own. In general, belief alignment can emerge as a natural consequence of each agent’s drive to minimize free energy within the right social context. This fosters synchronized behavior, shared narratives, and collective understanding, unifying individual agency, social influence, and collective behavior under the cohesive formalism of AIF [Friston et al., 2020, 2024]. However, as the brief analysis above reveals, such synchronization through belief alignment may come about as a result of the exertion of social power, highlighting the importance of understanding not just when this synchronization occurs, but also how it does so. As we conceptualized it, power is an agent’s capacity to shape and control the (semantic) “attractors” within a social or informational landscape. These attractors represent stable states or patterns that agents are drawn towards, which reduce uncertainty and create predictable social structures. Power enables agents to influence which states become salient or preferred for others, thereby controlling the flow of attention, shaping beliefs, and modulating the precision of predictions. The 20th and 21st century have shown, time and time again—especially so in the last 20 years—how mediatic domination (narrative control) results in the shaping of attractors which organize life.
@@
In leader-follower dynamics, a leader minimizes this divergence by exerting control over precision, amplifying others’ confidence in specific interpretations of their sensory data. For example, if the leader’s goal is clearly defined and they select social epistemic actions (more informative, albeit less efficient actions), they convey their intentions early and reduce uncertainty for followers [Maisto et al., 2024]. This approach encourages other agents to update their beliefs to align with the leader’s, effectively lowering KL divergence across the group. Even in cases where leadership is not predefined, an agent can assume a leadership role by adjusting precision to show high confidence in their beliefs or goals. This increase in precision draws other agents toward the confident agent’s stance, resulting in a convergence on shared beliefs. Precision-driven leadership can thus promote belief alignment without coercion, as agents gradually follow those who express strong, reliable predictions. In this way, the power to direct attention, structure narratives, and therefore modulate precision enables agents to actively shape collective beliefs, creating cohesion within groups and guiding group behavior through the minimization of belief divergence.
@@
The applications of this inevitably begin at reductive formalizations (and the math is never the territory [Andrews, 2021]), however, the implications of these arguments for complex social dynamics in the realms of, e.g., policy and governance are vast. Coordination dynamics at the level of simple perception-action cycles such as the ones analyzed in [Maisto et al., 2024], need to be framed in terms of their possibilities for novel understandings of social change processes: how should humanity—as one, as many—develop strategies with enough coherence and plasticity so as to resist the oppressive aspects of dominant narratives which do not benefit the coordination of the whole? What kinds of timeframes are we thinking about (in terms of generations, or the reorganization of our understanding of future time altogether)? Leader-based top-down control ‘works’ when established cooperation dynamics are already in place, and the same can be said about distributed control (either as presented by a socially-positive cybernetics (e.g., Beer) or “free market”, invisible hand narratives).
@@
A starting point for the analysis of such multi-agent interactions may be from the point of view of game theory. In particular, the concept of Free-Energy Equilibria (FEE) captures the notion of a joint policy in which all agents in a given interaction are minimizing their (Expected) Free Energy [Hyland et al., ]. Under certain conditions, it may be expected that interactions between agents tend to drive the system toward FEE. However, as is well understood in game theory, there are typically many equilibria of a game and depending on the starting point and interaction/learning dynamics, the system may converge to undesirable equilibria where a handful of agents exert significant control over others to their detriment. An important area of future study is thus to first establish a better understanding of what makes different equilibria more or less societally desirable, and then to understand how incentive landscapes may be shaped to navigate society toward better equilibria. Shared narratives, where agents are in the know about what is shared, are fundamental here, and this begins by paying attention to the (possible) rules in place, hence the need for formal methods. This is part of what critical metacognition entails: knowing how we are collectively navigating uncertainty, and taking account of already-given asymmetries. In the conclusion below, we explore a few salient implications of our arguments.